2. A general genetics class is asked to study three traits (all controlled by single genes with complete dominance) in a population of mice - coat color (black (w) white (w)), eye color (brown (b) - blue (b)) and tail shape (straight (k) - kinked (k)). Assuming the genes are linked use the following data to determine the linkage distance between the genes and the gene order. Be sure to correct for the double crossovers. Black Brown Straight - 452 Black Brown Kinked-67 Black Blue Straight - 3 Black Blue Kinked 33 White Brown Straight -31 White Brown Kinked 2 White Blue Straight -66 White Blue Kinked 453

//////////// Calculate the double crossover interference value //////////
Answer #1

Answer:

2. A general genetics class is asked to study three traits (all controlled by single genes with complete dominance) in a popu

Hint: Always recombinant genotypes are smaller than the non-recombinant genotypes.

Hence, the parental (non-recombinant) genotypes is w+b+k+ / wbk

1).

If single crossover occurs between w+&b+

Normal combination: w+b+/wb

After crossover: w+b / wb+

w+b progeny= 3+33=36

wb+ progeny = 31+2=33

Total this progeny =69

Total progeny = 1107

The recombination frequency between w+&b+ = (number of recombinants/Total progeny) 100

RF = (69/1107)100 = 6.23%

2).

If single crossover occurs between b+&k+..

Normal combination: b+k+/bk

After crossover: b+k/bk+

b+k progeny= 67+2=69

bk+ progeny = 66+3=69

Total this progeny = 138

The recombination frequency between b+ & k+ = (number of recombinants/Total progeny) 100

RF = (138/1107)100 = 12.47%

3).

If single crossover occurs between w+&k+..

Normal combination: w+k+/wk

After crossover: w+k/wk+

w+k progeny= 67+33=100

wk+ progeny = 66+31=97

Total this progeny = 197

The recombination frequency between w+ & k+ = (number of recombinants/Total progeny) 100

RF = (197/1107)100 = 17.80%

Recombination frequency (%) = Distance between the genes (cM)

w--------6.23cM-----b---------12.47cM--------------k

Expected double crossover frequency = (RF between w & b) * (RF between b & k)

= 6.23% * 12.47% = 0.0078

As heterozygote order of genes is w+b+k+ & wbk its genotype after double crossovers are w+bk+ & wb+k.

The observed double crossover frequency = 3+2/1107 = 0.0045

Coefficient of Coincidence (COC) = Observed double crossover frequency / Expected double crossover frequency

= 0.0045 / 0.0078

= 0.58

Interference = 1-COC

=1-0.58

=0.42